### v1

@@ 4.3.2 Offensive Talks
Figure 12: Similarity score of offensive talks mapped to MITRE over years
@@
Figure 13: Similarity score of offensive talks mapped to MITRE ATT&CK
@@
Figure 14: Cosine similarity score for the techniques within Execution.
@@
5 Limitations
The four main limitations of our study are: missing data (e.g. sponsorship for some conferences); processing speaker and company names; classifying abstracts against frameworks; and incomplete coverage of security conferences.
Missing Data Our comparative longitudinal study was impacted by changes in reporting and missing data. We could not collect some years of talks due to data unavailability. For example, the RSA conference was created in 1991 but we could only extract the program from 2017 onward. Similarly, there was a lot of missing sponsorship information, which led to a long list of exceptions to the sponsors column in Table 1.
@@ 5 Limitations
Real Names Counting authors can be complex because two unique names may correspond to the same individual (e.g. Emma Jones and Emma Q. Jones), and one name may correspond to multiple individuals (e.g. Emma Jones from New York and Emma Jones from California). Additionally, some authors may use pseudonyms rather than their real names. For simplicity, we counted unique names as unique people. We considered using affiliations to distinguish between two unique individuals with the same name, but researchers switch institutions too often for this to be reliable.
@@
Similar issues are present for company names. Company names are inconsistently abbreviated (Hewlett Packard vs HP), subsidiaries and even departments are listed (e.g. both Microsoft Corporation and Microsoft Research were listed as sponsors), and some companies change names (e.g. Facebook to Meta). We fixed such issues manually where we could, but we likely missed instances for less well known companies.
@@
Classification Mapping abstracts to categories from security frameworks is challenging because: (i) the task is inherently subjective; (ii) the categories are high-level with fuzzy boundaries; (iii) abstracts typically focus on specific technological systems; and (iv) even state-of-the-art LLMs are known to have errors. These factors help to explain why the agreement between the two NLP techniques (see Section 3.3) was disappointing. Ultimately, we focused on the LLM-based mapping because it is the state-of-the-art technique and it captures how language is used in real discussions in the LLM’s training set. This can be contrasted against the cosine similarity approach, which is trained on a bag-of-words extracted from official documents. However, it is not unreasonable to prefer a different mapping, such as using another NLP technique or another cyber security framework. To enable this, we open-source our data and analysis scripts.
@@
Conference Coverage Our comparative analysis (RQ3) would have been more interesting if we collected data from more conferences and also went further back in time. For example, the FIRST Conference on Computer Security Incident Handling would likely have covered more topics related to Respond, but from a more technical perspective than the insurance conferences. The core challenge to adding more data is that we could only semi-automated our scrapers in order to ensure data quality, and had to create a new parser for each conference. This made it costly to add additional conferences and also to scrape further back in time. Again, we hope that open-sourcing our data and analysis will allow future work to extend our work by adding new conferences.
6 Discussion
@@
We discuss the results and how our framing assumptions shaped the analysis.
@@
Speakers Academia displays more information sharing across conferences than industry, given many industry speakers present at just one conference and speakers rarely cross the insurance–InfoSec line (see Figure 6). It is unclear whether industry conferences would benefit from more cross-pollination. The call for broader information sharing is appealing, however it could be that more efficient communication is enabled by the shared language and assumptions within each community.
@@
A different question is whether sponsorship buys influence drezner2017ideas . Conferences with more sponsors have fewer words per abstract, which could be interpreted as sponsored conferences involving less complex talks. Even though Black Hat had the most sponsors, this appears to have little influence over who speaks. For example, the majority of sponsors (63%) never had an affiliated talk, and the majority (85%) of the speakers’ affiliated organizations never sponsored Black Hat. The main concern for academic conferences is that despite being privacy conferences, two of the five most common sponsors are companies that directly profit from behavioral advertising. The relationship between BigTech companies and researchers has been called into question by other researchers abdalla2021grey .
@@
A similar question is whether it is a problem that a small number of speakers/authors hold the majority of influence. For illustrative purposes, we can compare the Gini index for conferences (see Table 1) to the index of income inequality in specific countries. The distribution of talks across speakers is typically around 0.37–0.44, which is close to the post-tax distribution of income in relatively unequal societies like the US (0.375) and Mexico (0.42). All prestigious industry and academic talks have higher inequality than Germany (0.296). [^5] All figures sourced from the OECD using post-tax income: https://stats.oecd.org Inequality in academic talks is not necessarily a cause for concern. Scientific conferences should provide a platform for the ‘best’ science and that likely means the ‘best’ scientists will present more often. However, an unequal distribution can also result from a narrow perspective of what constitutes ‘good’ research. It is worth noting that ACM CCS was the most unequal academic conference across both approaches to quantifying the Gini Index.
Topics One way to improve inequality metrics is to expand the topics that are presented at top tier academic security conferences, allowing new researchers to present ideas and findings. For example, there is comparatively little research about the Identify function that concerns security management. A first step would be to understand how security is managed at present hielscher2023employees , which can motivate technical solutions to management problems. For example, Asset Management can be inferred from network traffic dekoven2019measuring ; Risk Assessment can be framed as ML prediction liu2015cloudy ; and Governance can involve technical audits rahaman2019security ; mahmud2020cardpliance .
@@
Turning to offensive talks, the academic and industry talks are much more evenly distributed across the techniques in MITRE ATT&CK. The one exception is information gathering that occurs pre-compromise, which is rarely discussed at conferences. Preventing resource development is perhaps better conducted by governments, such as by disrupting criminal markets for exploits allodi2015then . The lack of offensive talks at insurance conferences raises the question of whether there is a sufficient understanding of attacker operations. One potential reason this does not happen is insurers rely on lawyers to manage incidents, and those lawyers do not report findings due to confidentiality concerns woods2023lessons .
@@
Variance The insurance conferences were the biggest outliers because they focus on management and resilience more than technological controls and attacker techniques. This confirms the finding that insurers are more focused on post-incident response than pre-incident mitigation wolff2022cyberinsurance . For the MITRE ATT&CK framework mapping, the only offensive conference (DEFCON) is comparatively focused on how compromise was achieved (execution, persistence and initial access), while the academic conferences focus on activities that take place after compromise.
@@
Turning to changes over time, there is an increasing focus on the Identify function and perhaps a weakening focus on Protective Technology and Detect functions. This contradicts the prediction of Schneier that the 2010s would be the decade of response schneier2014future . Instead, there appears to be a subtle shift to security management. However, we should not overstate these effects as the ‘declining’ categories remain the most frequently discussed in absolute terms.
Assumptions In mapping abstracts to cybersecurity frameworks, we imply that there should be equal coverage across topics of the framework. Our core assumption is that more speakers/authors and different topics is better. We justify this by pointing to the benefits researchers gain from learning about areas outside of their specialism. On the other hand, focusing on ever more specific areas of security facilitates deeper understanding and more progress. This is an example of the explore vs exploit trade-off with an unknown objective function, namely maximizing global security knowledge. We acknowledge that reasonable readers may disagree with our emphasis on explore.
@@
Future Work We focused on speakers/authors but did not consider listeners/readers. Doing so is crucial to understanding the impact of conference presentations. Collecting this information risks introducing some form of surveillance that undermines the privacy of attendees. One privacy-preserving approach is to scrape the number of views that uploaded video recordings received and to study the comments. Alternatively, an (imperfect) academic proxy for impact on readers is citation count, but the industry equivalent requires more sophisticated research designs.
@@ 7 Conclusion
The results show academics moved freely between the big four venues, but just 5% of academic authors also presented at InfoSec or Insurance conferences. The equivalent figure was 4% for insurance speakers, and 7% for InfoSec speakers, which means there is scope for presenters to move between disciplines. We also showed there is an unequal distribution of talks across speakers, with this inequality varying across conferences. In terms of sponsors, academic conference are dominated by BigTech firms, whereas InfoSec conference sponsors are predominantly cybersecurity vendors.
@@
All conferences contained more offensive than defensive talks, with the exception of DEFCON. Offensive abstracts are distributed evenly across the categories of MITRE ATT&CK, apart from Reconnaissance and Resource Development. Defensive abstracts focused on Protective Technology and categories related to monitoring. Academic conferences tend to overlook the Identify and Recover functions from the NIST Cybersecurity Framework.
@@
These findings support a more nuanced understanding of security information sharing, in which success should be evaluated with regards to specific channels and institutions. Researchers have been disappointed by the quality of information shared via online sources acar2016you ; fischer2017stack and threat intelligence feeds li2019reading ; bouwman2020different . However, there is more success when it comes to vulnerability management. Bug bounties incentivize the discovery of new information finifter2013empirical ; sridhar2021hacking . Further, information about fixes is widely available, the problem lies in acting on the information by installing patches frei2006large ; li2017large . Our results suggest security conferences should be considered a success story, closer to patch management than threat intelligence sharing. This is perhaps not surprising given attendees, sponsors and research founders invest considerable resources in attending and sponsoring these conferences.
Going forward, organizers of prestigious conferences should ensure there is a range of speakers and topics. DEFCON is perhaps a good example in that it refuses sponsorship, evenly covered all but two areas of the MITRE ATT&CK framework, and was one of the most equal conference in terms of the distribution of talks across speakers. An interesting follow-on study would speak to conference organizers to enrich our quantitative analysis with qualitative insights about the unobservable aspects of conference organization.

### v2

@@ 4.3.2. Offensive Talks
Figure 12. Similarity score of offensive talks mapped to the MITRE ATT&CK framework over time (2014–2022).
@@
Figure 13. Similarity score of offensive talks mapped to the MITRE ATT&CK framework across conferences.
@@
Figure 14. Cosine similarity score for talks classified under the Execution technique category of MITRE ATT&CK framework across conferences.
@@
5. Threats to Validity and Limitations
Our study has several limitations that affect the interpretation and generalizability of the results.
Threats to External Validity (Generalizability) Missing Data: Our comparative longitudinal study was impacted by changes in reporting and missing data. We could not collect some years of talks due to data unavailability.
Our longitudinal dataset is affected by occasional missing years or incomplete sponsor information, particularly for some industry conferences. For example, the RSA conference was created in 1991 but we could only extract the program from 2017 onward. Similarly, there was a lot of missing sponsorship information, which led to a long list of exceptions to the sponsors column in Table 2. These omissions mean that 13 of the 74 expected conference-year sponsorship data points (approximately 17.6%) are missing from the analysis. This lack of data systematically biases aggregate statistics toward the better-documented conferences and limits the completeness of the longitudinal trend analysis. Future releases of our dataset will incorporate missing years as additional data become publicly available.
@@ 5. Threats to Validity and Limitations
Threats to Construct Validity (Measurement) Real Names: Counting authors can be complex because two unique names may correspond to the same individual (e.g. Emma Jones and Emma Q. Jones), and one name may correspond to multiple individuals (e.g. Emma Jones from New York and Emma Jones from California). Additionally, some authors may use pseudonyms rather than their real names. For simplicity, we counted unique names as unique people. We considered using affiliations to distinguish between two unique individuals with the same name, but researchers switch institutions too often for this to be reliable.
@@
Similar issues are present for company names. Company names are inconsistently abbreviated (Hewlett Packard vs HP), subsidiaries and even departments are listed (e.g. both Microsoft Corporation and Microsoft Research were listed as sponsors), and some companies change names (e.g. Facebook to Meta). We fixed such issues manually where we could, but we likely missed instances for less well known companies. This fragmentation is estimated to affect less than 2% of total unique entities after manual correction, primarily impacting measures of speaker concentration and sponsor overlap but not materially changing overall long-term trends.
@@
Classification: Mapping abstracts to categories from security frameworks is challenging because: (i) the task is inherently subjective; (ii) the categories are high-level with fuzzy boundaries; (iii) abstracts typically focus on specific technological systems; and (iv) even state-of-the-art LLMs are known to have errors, let alone the version we used that is now far from state-of-the-art. These factors help to explain why the agreement between the two NLP techniques (see Section 3.3) was disappointing. Ultimately, we focused on the LLM-based mapping because it was the best available technique at the time of the initial analysis and it captures how language is used in real discussions in the LLM’s training set. This was also confirmed by the results of manual annotation: 264 abstracts (2.7% of the corpus) were validated, achieving a 92.4% agreement (95% CI ±3.1 pp) and a Cohen’s kappa coefficient of 0.89 with human annotators. However, residual errors and category ambiguities are still possible. This can be contrasted against the cosine similarity approach, which is trained on a bag-of-words extracted from official documents. However, it is not unreasonable to prefer a different mapping, such as using another NLP technique or another cybersecurity framework. To enable this, we open-source our data and analysis scripts (Walter et al., 2024).
@@
Conference Coverage: Our comparative analysis (RQ3) would have been more interesting if we collected data from more conferences and also went further back in time. For example, the FIRST Conference on Computer Security Incident Handling would likely have covered more topics related to Respond, but from a more technical perspective than the insurance conferences. The core challenge to adding more data is that we could only semi-automated our scrapers in order to ensure data quality, and had to create a new parser for each conference. This made it costly to add additional conferences and also to scrape further back in time. Again, we hope that open-sourcing our data and analysis will allow future work to extend our work by adding new conferences.
Threats to Internal Validity (Causality): Our results are descriptive, not causal. While we discuss possible explanations for observed patterns (e.g., topic prevalence or speaker concentration), these should not be interpreted as evidence of causation. Since our study is based on descriptive data, we explicitly avoid making causal claims to prevent threats to internal validity.
6. Discussion
@@
This section discusses the results and how our framing assumptions influenced the analysis.
@@
Speakers Academic conferences display greater information sharing across venues than industry. Many industry speakers present at only one conference, and speakers rarely cross the insurance–InfoSec line (see Figure 6). It is unclear whether industry conferences would benefit from more cross-pollination. The call for broader information sharing is appealing, however it could be that more efficient communication is enabled by the shared language and assumptions within each community.
@@
A common question is whether sponsorship buys influence (Drezner, 2017). Conferences with more sponsors have fewer words per abstract, which could be interpreted as sponsored conferences involving less complex talks. However, the relationship between sponsorship and speaker selection appears limited. For example, at Black Hat — the event with the largest number of sponsors — 63% of sponsors never had an affiliated talk, and 85% of speakers’ organizations never sponsored the conference. Nevertheless, sponsorship patterns raise concerns for academic conferences. Despite focusing on privacy, two of the five most frequent sponsors of academic conferences directly profit from behavioral advertising. The relationship between BigTech companies and researchers has been called into question by other researchers (Abdalla and Abdalla, 2021).
@@
A similar question is whether it is a problem that a small number of speakers/authors hold the majority of influence. For illustrative purposes, we can compare the Gini index for conferences (see Table 3) to the index of income inequality in specific countries. The distribution of talks across speakers is typically around 0.37–0.44, which is close to the post-tax distribution of income in relatively unequal societies like the US (0.375) and Mexico (0.42). All prestigious industry and academic talks have higher inequality than Germany (0.296). [^6] All figures sourced from the OECD using post-tax income: https://stats.oecd.org (accessed 19 October 2025) Inequality in academic talks is not necessarily a cause for concern. Scientific conferences should provide a platform for the ‘best’ science, and that likely means the ‘best’ scientists will present more often. However, an unequal distribution can also result from a narrow perspective of what constitutes ‘high-quality’ research. Among academic venues, ACM CCS displayed the highest inequality across both Gini computation approaches.
Topics Expanding the topical diversity of academic security conferences could mitigate inequality by allowing new researchers to present alternative perspectives. Some areas - particularly the Identify function within the NIST Cybersecurity Framework - remain underrepresented. Understanding current security management practices (Hielscher et al., 2023) could inspire technical approaches to management problems. For instance, Asset Management can be inferred from network traffic (DeKoven et al., 2019); Risk Assessment can be framed as ML prediction (Liu et al., 2015); and Governance can involve technical audits (Rahaman et al., 2019; Mahmud et al., 2020).
@@
Turning to offensive talks, their distribution is much more even across the techniques in MITRE ATT&CK, spanning both academic and industry conferences. The one exception is information gathering that occurs pre-compromise, which is rarely discussed at conferences. Preventing resource development is perhaps better conducted by governments, such as by disrupting criminal markets for exploits (Allodi et al., 2015). The lack of offensive talks at insurance conferences raises the question of whether there is a sufficient understanding of attacker operations.
@@
Variance Insurance-focused events are clear outliers: they emphasize management and resilience rather than technical controls or attacker techniques. This confirms the finding that insurers are more focused on post-incident response than pre-incident mitigation (Wolff, 2022). Within MITRE ATT&CK mappings, the only offensive conference (DEFCON) is comparatively focused on how compromise was achieved (execution, persistence and initial access), while the academic conferences focus on activities that take place after compromise.
@@
Over time, there appears to be a gradual shift towards the Identify function and away from Protective Technology and Detect functions. This trend contradicts the prediction of Schneier that the 2010s would be the decade of response (Schneier, 2014). Instead, there appears to be a subtle shift to security management. However, the apparent declines should not be overstated, as the ‘declining’ categories remain the most frequently discussed in absolute terms.
Assumptions In mapping abstracts to cybersecurity frameworks, we imply that there should be balanced coverage across framework categories. Our core assumption is that a broader range of speakers and topics implies a healthier research ecosystem. While diversity has known benefits, specialization also drives depth and innovation. This tension reflects the classic explore–exploit trade-off, where the objective — maximizing global security knowledge — remains undefined. We acknowledge that reasonable readers may disagree with our emphasis on explore.
@@
Future Work We focused on speakers but did not consider listeners/readers. Doing so is crucial to understanding the impact of conference presentations. Collecting this information risks introducing some form of surveillance that undermines the privacy of attendees. One privacy-preserving approach is to scrape the number of views that uploaded video recordings received and to study the comments. Alternatively, an (imperfect) academic proxy for impact on readers is citation count, but the industry equivalent requires more sophisticated research designs.
@@ 7. Conclusion
The results show academics moved freely between the big four venues, but just 5% of academic authors also presented at InfoSec or Insurance conferences. The equivalent figure was 4% for insurance speakers, and 7% for InfoSec speakers. This suggests there is limited information sharing across academia, InfoSec and insurance; a potential area for improvement. We also showed there is an unequal distribution of talks across speakers, with this inequality varying across conferences. In terms of sponsors, academic conference are dominated by BigTech firms, whereas InfoSec conference sponsors are predominantly cybersecurity vendors. Unequal distribution of speaking time or sponsorship sources can be a sign that parties have the power to influence which security topics are studied and which solutions are adopted.
@@
Defensive talks were more common than offensive and neutral ones at all conferences, except for DEFCON, where offensive topics dominated due to its hacker-focused nature. This allows us to quantify the balance of offensive and defensive papers at academic conferences, which have been criticized for focusing on attack papers (Herley and Van Oorschot, 2017). Offensive abstracts are distributed evenly across the categories of MITRE ATT&CK, apart from Reconnaissance and Resource Development. Defensive abstracts focused on Protective Technology and categories related to monitoring. Academic conferences are less likely to focus on the Identify and Recover functions from the NIST Cybersecurity Framework.
@@
These findings support a more nuanced understanding of security information sharing, in which success should be evaluated with regards to specific channels and institutions. Researchers have been disappointed by the quality of information shared via online sources (Acar et al., 2016; Fischer et al., 2017) and threat intelligence feeds (Li et al., 2019b; Bouwman et al., 2020, 2025). However, there is more success when it comes to vulnerability management. Bug bounties incentivize the discovery of new information (Finifter et al., 2013; Sridhar and Ng, 2021). Further, information about fixes is widely available, the problem lies in acting on the information by installing patches (Frei et al., 2006; Li and Paxson, 2017). Our results suggest security conferences should be considered a success story, closer to patch management than threat intelligence sharing. This is perhaps not surprising given attendees, sponsors and research founders invest considerable resources in attending and sponsoring these conferences. An interesting follow-on study would speak to conference organizers to enrich our quantitative analysis with qualitative insights about the unobservable aspects of conference organization.
Going forward, organizers of prestigious conferences should ensure there is a range of speakers and topics. To increase diversity of speakers, program committees could explore proposals like: (i) reserving slots for authors/institutions who have not presented before; and (ii) introducing a maximum number of submissions per individual (implemented by USENIX Security, for example). To increase topic diversity, conferences could edit the call for papers to address under-studied topics according to our results, and possibly lower the bar for acceptance for these topics—a policy not unlike special issues for journals. We only raise these policies as provocations, acknowledging such policies will be argued to be in tension with the principle of meritocracy. Nevertheless, we hope our article can prompt discussions among program committees.
Appendix: Supplementary Figures and Tables
Figure 15. Mean similarity score (averaged across GPT-based and cosine similarity measures) for defensive talks mapped to the NIST Cybersecurity Framework.
Figure 16. Mean similarity score (averaged across GPT-based and cosine similarity measures) for offensive talks mapped to the MITRE ATT&CK Framework.
Figure 17. Similarity score of Black Hat defensive talks mapped to the NIST Cybersecurity Framework from 2014 to 2022.
Figure 18. Similarity score of IEEE S&P defensive talks mapped to the NIST Cybersecurity Framework from 2014 to 2022.
Figure 19. Top sponsors of InfoSec and cyber insurance conferences aggregated over 2014–2022. Numbers above the bars indicate each sponsor’s mean sponsorship tier. Note: sponsor data for some InfoSec conferences were incomplete.
Figure 20. Similarity score of USENIX Security defensive talks mapped to the NIST Cybersecurity Framework from 2014 to 2022.
Figure 21. Similarity score of Black Hat offensive talks mapped to the MITRE ATT&CK Framework from 2014 to 2022.
Figure 22. Similarity score of DEFCON offensive talks mapped to the MITRE ATT&CK Framework from 2014 to 2022.
Figure 23. Similarity score of IEEE S&P offensive talks mapped to the MITRE ATT&CK Framework from 2014 to 2022.
Figure 24. Similarity score of USENIX Security offensive talks mapped to the MITRE ATT&CK Framework from 2014 to 2022.